Machine hours and electricity costs for Wells Industries for last year were as follows:

Machine

Electricity

Month

Hours

Costs

January

2,000

$ 9,200

February

2,320

10,500

March

1,520

6,750

April

2,480

11,500

May

3,040

14,125

June

2,640

11,000

July

3,280

12,375

August

2,800

11,375

September

1,600

7,750

October

2,960

13,000

November

3,760

15,500

December

3,360

13,875

Required: Show your Calculations!

A. Using the high-low method, develop an estimate of variable electricity costs per machine hour.
B. Using the high-low method, develop an estimate of fixed electricity costs per month.
C. Using the high-low method, develop a cost function for monthly electricity costs.
D. Estimate electricity costs for a month in which 3,000 machine hours are worked.

Question 41 options:
Answer #1
Machine Electicity
hours cost
high 3,760 15,500
low 1,520 6,750
Difference 2,240 8,750
A Variable cost per MH =8750/2240
3.91
B Fixed electicity costs per month
15,500 - (3760*3.91)
798.4
C Cost Function
y                = a + bx
total cost (y) = 798.40 + 3.91x
total electricity cost = 798.40 +3.91 machine hours
d Estimated Electricity cost
798.40 +3.91*3000
12528.4 answer